OpenStudy (anonymous):

(b) A related rule of thumb is that a typical man who is 70 inches tall weighs 170 pounds. On the basis of these two rules of thumb, use a formula to express the trend giving weight w as a linear function of height h. w(h)= ?

jimthompson5910 (jim_thompson5910):

it sounds like weight (w) is directly related to height (h) so we have direct variation w = k*h where k is some constant (a fixed number)

jimthompson5910 (jim_thompson5910):

plug in the given weight and height, then solve for k w = k*h 170 = k*70 I'll let you isolate k

OpenStudy (anonymous):

k=2.4

jimthompson5910 (jim_thompson5910):

I'm getting 170/70 = 2.42857142857142 if you round that to one decimal place, then you get k = 2.4 so it looks good

jimthompson5910 (jim_thompson5910):

so `w(h) = 2.4*h` whatever the height is, the weight will be approximately 2.4 times the height weight = 2.4 times height

OpenStudy (anonymous):

but w= 2.4 * 70 isn't it, its slope

jimthompson5910 (jim_thompson5910):

well they want a general formula not just that specific value

jimthompson5910 (jim_thompson5910):

so you replace the 70 with some general height h

jimthompson5910 (jim_thompson5910):

to get `w(h) = 2.4*h`

OpenStudy (anonymous):

that isn't it either :(

OpenStudy (anonymous):

its 5h-180 thanks!

jimthompson5910 (jim_thompson5910):

hmm strange, it seems like they're converting but I'm not sure

jimthompson5910 (jim_thompson5910):

does it use another value other than (70, 170)?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

I notice it says part B what was part A about?